OpenStudy (anonymous):

derivative of f(x)= x^pi^2+(pi^2)^x

OpenStudy (anonymous):

Remember that \[\pi ^{2}\] is just a constant. What is the derivative of:\[f(x) = x ^{c} + c ^{x}\] \[f'(x) = \ c* x ^{(c-1)} + \ln(c) {c} ^{x}\] \[f'(x) = \pi ^{2} x ^{(\pi ^{2}-1)} + \ln(\pi ^{2}) {\pi ^{2}} ^{x}\]